Caliber Corp. currently pays no dividends because it requires its internally generated funds to be used to fund a research intensive project over the next three years. The project is expected to be completed after this time and by the end of the fourth year, the company will pay its first dividend of $2.00. The dividends will be expected to grow at the rate of 20% over the following five years. After this time period, dividends will grow at the rate of 2% into the foreseeable future. Assuming you require a 15% rate of return, compute the value of this company's stock.
Answer #1

Dividend in year 4 =2
D5 =2*1.20
D6 =2*1.20^2
D7=2*1.20^3
D8=2*1.20^4
D9=2*1.20^5
D10 =2*1.20^5*(1+2%)
Terminal Value at year 9 =2*1.20^5*(1+2%)/(15%-2%) =39.0475
Price of Stock =2/1.15^4+2*1.20/1.15^5+2*1.20^2/1.15^6+2*1.20^3/1.15^7+2*1.20^4/1.15^8+2*1.20^5/1.15^9+39.0475/1.15^9
=18.75